OpenStudy (anonymous):

Medal given! Question 1:Solve the following system of equations by using the multiplication method: 2x - 7y = 9 3x - 4y = -6 Question 2:Solve the following system of equations using multiplication method: 5x + 3y = 1 x + 2y = 10. Show your work

OpenStudy (aaronq):

The purpose of multiplying the equations is that you end up with the same variable with the same coefficient so when you add them they cancel out and you're left with only one. I'll do the first one: (1) 2x - 7y = 9 (2) 3x - 4y = -6 Multiply (1) by -3 and (2) by 2 (1) -6x + 21y = -27 (2) 6x - 8y = -18 Now add them: -6x+6x + 21y -8y= -27 + ( -18) 13y=-45 y= \(\dfrac{-45}{13}\) You can find x by pluggin this into the prior equation 6x - 8( \(\dfrac{-45}{13}\)) = -18 \(x=\dfrac{21}{13}\)

OpenStudy (aaronq):

oh wait i made a mistake. (1) -6x + 21y = -27 (2) 6x - 8y = \(\color{red}{-12}\)

OpenStudy (aaronq):

so, \(y=-3\) -6x + 21y = -27 -6x + 21(-3) = -27 \(x=-6\)

OpenStudy (anonymous):

That actually helped me out a lot.. Thank you, I understand how to do those now @aaronq

OpenStudy (anonymous):

Can you help me with another? @aaronq

OpenStudy (aaronq):

No problem, glad i could help! sure post away

OpenStudy (anonymous):

Find - 3x2y(6x2y2 - 8x +12y) Choose one answer. a. -9 x3y3 - 8x3y +12x2y2 b. - 18x4y3 + 24x3y + 36x2y2 c. - 9x4y3 + 11x3 + 36x2y d. - 18x2y3 -11x3y - 15x2y2

OpenStudy (aaronq):

is it this? \( - 3x^2y(6x^2y^2 - 8x +12y)\)

OpenStudy (aaronq):

i mean the original function with the exponents

OpenStudy (anonymous):

Yess

OpenStudy (aaronq):

k, so you pretty much have to distribute the \( - 3x^2y\) at the front. you can do it one at a time or all at once. I'll do it one at a time first. \( - 3x^2y(6x^2y^2 - 8x +12y)\) \( x^2y[(-3)*6x^2y^2 - (-3)*8x +(-3)*12y]\) \( x^2y[-18x^2y^2 +24x -36y]\) now the x^2 \( y[-18x^4y^2 +24x^3 -36yx^2]\) now the y \( -18x^4y^3 +24x^3y -36x^2y^2\)

OpenStudy (anonymous):

Oman.. I feel really dumb. That's like the easiest thing to do in this whole class ahhaha. Thanks for breaking it down ;p

OpenStudy (aaronq):

haha dont feel dumb. You should do these things on actual paper, it's much much easier to see as long as you keep things neat. No problem though !